# Demystifying Platform Requirements for Diverse LLM Inference Use Cases

### v1

@@ III-D Metrics for LLM Serving
IV GenZ : Generative LLM Analyzer
In this work, we aim to understand the platform-wide requirements to meet the Service Level Objectives (SLOs) of LLMs for different use cases. Platforms are defined as collections of various accelerators connected through an interconnect network, such as Nvidia HGX[26], Google TPU pods[8], Intel HLS-Gaudi2[17], AMD Instinct MI300X Platform [3], and many more [12, 114][21, 2, 40].
There is a lack of an existing tool/framework in the community that can help us study LLM inference at various operating points with different inputs. There exist several Neural Processing Unit (NPU) compute simulators [116, 94, 107] but are focused on modeling the compute behavior in detail (e.g., dataflow) and lack support for modeling distributed platforms. Meanwhile, simulators like ASTRA-sim [110] focus on communication optimizations for distributed training and do not model the specific prefill-decode pipeline required for distributed inference. To address the issue, we build a first-order analytical tool, GenZ [^2] https://github.com/abhibambhaniya/GenZ-LLM-Analyzer. GenZ has three key components: 1) Model Profiler, 2) NPU Characterizer, and 3) Platform Characterizer. We show an overview of GenZ in Fig. 2 and discuss each of its components in the following sections.
IV-A Target Use Cases
A key difference among LLMs compared to previous DNN models (such as ResNet [74]), is that the model execution shapes vary widely as the use case changes. A use case is characterized by three key features :
  1. 1.
    Beam Size, $S_{b}$: Number of parallel outputs maintained during the decode phase. Typically value: 2-4.
  2. 2.
    Number of Input Tokens, $\tau_{p}$: Number of tokens that the user has passed as the input prompt. This value varies widely with different applications. $T_{TTFT}$ depends primarily on $\tau_{p}$.
  3. 3.
    Number of Output Tokens, $\tau_{d}$: Number of tokens to be generated during the decode phase in an auto-regressive fashion. Similar to $\tau_{p}$, this value varies widely with different applications.
Table IV shows the five use cases we considered for this work. These have varying input token lengths, output token lengths, and latency constraints. These use cases represent different real-world applications, such as chat Services, code generation, summarization, and question-answering. Additionally, Retrieval augmented generation (RAG)[97, 72] is used in almost all question-answering-based LLM agents. We also model this as a separate use case where the input tokens have context appended to the input prompt.
We use the term LLM workload to refer to the specific job running on the hardware platform and placing specific computational demands, and define it as follows:
$Workload=\{Usecase,LLMModel,Optimizations\}$ where
$Optimizations=\{BatchSize,Precision,Parallelism\}$
Real Life Application
@@ IV-A Target Use Cases
Beam Width
TTFT
TPOT
Question Answering 1000/200 4 0.2 10
Chat Services 3000/1000 2 0.2 10
QA + RAG 10000/200 4 0.4 10
Text Summarization 15000/1000 4 2 20
Code Generation 20000/50 4 0.5 20
TABLE IV: Representative Use Cases of LLM models and their representative input hyperparameters.
We use various hugging face datasets[14] along with LLaMA tokenizer[122] to estimate input and output tokens for various use cases. We use dolphin dataset[4] for chat services token estimation. We use LongBench[46] for average token estimation for text summarization [32, 34, 35, 37] and code generation[33, 36]. For RAG-based applications, the context length can vary from a few thousand tokens to tens of thousands of tokens. We use MLDR[55], a Multilingual Long-Document Retrieval dataset built on Wikipedia, to estimate additional context tokens to user input.
IV-B Model Profiler
@@
Table I shows the details of various LLMs that we study in this work. We consider five different models that serve as a representative set of current state-of-the-art LLMs. We choose models representing the complete spectrum of model size ranging from 7B to 1.8T parameter LLMs. We also have models with different architectures, including MoE-based LLMs (Mixtral-8x7B and GPT4). LLaMA2, LLaMA3, Mixtral, and GPT3 architecture are available openly. We represent GPT4 as a 1.8T parameter mixture-of-experts model, with 120 layers[41, 9, 10]. A single layer of GPT4 has 16 experts of  [^1] B parameters each, and two experts are activated for each token. There are 55B shared parameters for attention. Each forward pass inference (generation of 1 token) utilizes about 280B parameters and about 560 TFLOPs. A dense 1.8 trillion parameter model would need about 3,700 TFLOP for each inference. We also develop a futuristic Super-LLM model using a scaled-up MoE model, super LLM. There are 32 experts per layer and about 10T parameters in total. It is a 128-layer deep model with 4 expert activations per token.
@@
For each new model, GenZ uses the huggingface AutoModels[16] to determine the exact shape of each operator in the model. Since all layers in the LLM are identical, we only need to save the operator dimension of one layer. All LLMs’ layer dimensions can be modeled using a few variables, as shown in Table I. Using these stored variables, we calculate the TFLOPS, weight size, KV Cache Estimation, and metrics like arithmetic intensity.
@@
Storing these model operators offline allows us to profile larger context lengths for any LLM model quickly. For example, LLaMA2-7B[123] has a context length limit of 4K[123], but using this offline modeling, we can extend the context length to any size.
@@
We also model popular optimizations in this stage of model profiling, including Kernel Fusion (Flash Attention[61, 60], FLAT[87], Segment KV Caching[128]) and model quantization (FP16, INT8/FP8, INT4/FP4). Since this work focuses on LLM inference, we use INT8 model quantization for all our experiments and results unless specified otherwise.
When profiling MoE models on GenZ, for decode, we assume that irrespective of batch size, only ‘k’ experts are selected at each layer of each decoding step. In Table I, we indicate the value of ‘k’ as the experts selected in the last column. For Prefill, we assume that all experts would be activated due to the large number of tokens, irrespective of batch size.
IV-C NPU Characterizer
Our smallest hardware unit is the accelerator (alternatively referred to as NPU), as shown in Fig. 2. We assume each NPU has a certain amount of efficient compute cores, which can perform $FLOPS$ number of operations per second. In an effort to keep the design space limited, we assume the on-chip memory is enough to achieve maximum computing efficiency. Additionally, we incorporate a computing efficiency factor to account for inefficiencies caused by software and memory synchronization issues.
Each NPU provides access to two external memories (fast and slow). Faster (smaller) memory could be something like an HBM/DDR Memory Bank, while Slower (Larger) memory could be PCIe-accessible CPU memory or CXL-accessible SSD/Flash. Most model weights and activations would prioritize using the Faster memory, however, if the capacity is not enough, data could be potentially offloaded to slower memory. While this would drastically reduce the inference performance, it would still enable inference rather than issuing an out-of-memory (OOM) error.
For the system, we analyze the model performance on an operator-by-operator basis. We follow a roofline-based approach to calculate each operator’s runtime on the accelerator.
$T_{mem}$ $=max(\frac{min(M_{op},S_{fast})}{BW_{fast}\times Eff_{fmem}},\frac{M_{op}-S_{fast}}{BW_{slow}\times Eff_{smem}})$ (1)
$T_{op}$ $=max(\frac{C_{op}}{FLOPS\times Eff_{C}},T_{mem})$ (2)
$S_{fast}$ is a partition of fast memory allocated to each operator. If $M_{op}>S_{fast}$, the additional data of the operator would be offloaded to the slower memory. Since all parameters, model weights, and KV cache are used uniformly, a fixed portion of the parameters will be offloaded. Using an ideal prefetching scheduler, each operator’s time to read from slower memory could be overlapped with the memory fetches from high-speed memory. Thus, total memory read time would be as shown in Equation 1.
IV-D Platform Characterizer
We define an inference platform as multiple NPUs connected using an interconnection network (ICN) as shown in Fig. 2. An ICN can have different topological connections, link latencies and varying link bandwidth. Typically distributed LLMs are served using five different types of parallelism strategies, Data Parallel (DP), Tensor parallel (TP) [120], Pipeline parallel (PP) [78], Expert parallel (EP)[65] (Only for MoE models) and Sequence parallel (SP)[81, 99] (primarily for training long sequence).
For TP LLM, an all-reduce communication primitive is required twice per layer. Thus a total of two×$\times$ number of layers all-reduce calls would be made per forward pass of the LLM.
We model all-reduce communication as explained in [44, 110, 127]. For all-reduce of message size M, where M is each accelerator’s initial/final data size, and N is the number of accelerators among which the model is distributed in a tensor parallel fashion. Each accelerator sends M/N-sized data to N-1 neighboring accelerators. All-Reduces time is modeled as:
$T_{AR}$ $=T_{Warmup}+2(N-1)(T_{Link}+\frac{M/N}{BW_{link}Eff_{link}})$ (3)
$T_{N2N}$ $=T_{Warmup}+T_{Link}+\frac{M}{BW_{link}Eff_{link}}$ (4)
For PP distribution of LLMs, only node-to-node message passing is required. In this case, the time to transfer a message of size M is defined as $T_{N2N}$.
For EP, two all-to-all would be required among all the nodes, which share experts of a layer. SP requires all gather and reduce scatter per layer.
In our modeling, we have a knob to decide whether to overlap collectives with compute tasks or execute them sequentially. Coordinating the simultaneous execution of collective communication and operator computation is an open research challenge [108]. Since existing SOTA LLM inference serving frameworks [95, 105, 28] do not incorporate these techniques, we opted to model the interconnect collective communication as non-overlapping. Considering everything, model execution time is defined by (6).
$T_{layer}$ $=\sum_{Ops}T_{op}+2T_{AR}$ (5)
$T_{model}$ $=Layers\times T_{layer}+(PP-1)\times T_{N2N}$ (6)
Figure 4: Comparing GenZ LLM inference modeling against vLLM inference. We vary the input batch size for LLaMA2-7B and LLaMA2-13B on H100 with different input prompt sizes and different LLM models. Different lines represent different input prompt sizes in each graph.
IV-E Validation Studies for GenZ
@@
We validate GenZ on a real hardware platform: A HGX[26] box with 8 NVIDIA H100 SXM GPUs (80 GB) fully connected by NVLinks. We conduct two types of studies: (i) Trends of prefill and decode on a single accelerator system to validate the accelerator modeling by GenZ. (ii) Collective times with different platform sizes to validate the platform interconnect modeling.
Single accelerator platform validation: For the first study, we serve LLaMA2-7B and LLaMA2-13B models using vLLM[96] framework on a H100 GPU. For each model, we run 2 use-cases ($S_{b}=1,\tau_{p}=544,\tau_{d}=200$) and ($S_{b}=1,\tau_{p}=2173,\tau_{d}=200$) and get LLM serving metrics (TTFT, TPOT). We compare this to profiling results generated by GenZ . Since the aim of GenZ is to give first-order analysis, we plug-in empirically measured efficiency factors of 60% for the GPU Tensor TFLOPs and 70% for GPU memory bandwidth. Fig. 4 shows both models’ TTFT and TPOT time comparisons. There is a 5.88% and 3.86% geomean error for TTFT and TPOT between real values and modeled values. This confirms that the trends produced by GenZ align with those observed on actual hardware.
@@
Platform interconnect validation: To verify the collectives time at platform scale, we benchmark all-reduce communication primitive with nccl-tests[24], a communication primitive performance benchmark for NVIDIA GPUs. GenZ ’s collective communication times are validated with a platform size of 2 GPUs, 4 GPUs, and 8 GPUs. For collective communications, we observed an average link efficiency of 75% for NVLINK, which gives an effective link BW of 350 GB/s for each GPU in HGX box. We profile all combinations of use case (Table IV) and model (Table I and collect their all-reduce (AR) message sizes. The message size of each AR call is very small ($<$ 128 KB) for decode, while for prefill, per call message size is a few hundred MBs. Fig. 5 compares the real hardware latency for three different platform sizes and the corresponding latency generated by GenZ . Since the prefill and decode stages have stark differences in message size, we verify the collective time for prefill and decode independently. We found that for decode message sizes, the link latency, $T_{Link}$ is the dominating, thus the latency seems almost constant. While prefill AR messages are large enough, that the link bandwidth, $BW_{Link}$, is the main contributor to collective time. Collective for all datapoints, there is a 3.89% and 2.7% geomean error for decode and prefill message sizes between real values and GenZ values.
(a) Decode Message sizes
(b) Prefill Message size
@@
V Workload Optimization Analysis with GenZ
For LLM inference, batch size, precision, and parallelism are considered optimizations aimed at improving performance, efficiency, and resource utilization without modifying the core implementation of the LLM. It’s important to note that while these optimizations can improve inference performance, some of them may also introduce trade-offs or limitations. For example, lower precision may result in some loss of accuracy, while parallelism strategies can introduce communication overhead or require specialized hardware configurations. In this section, we study the effect of input batching and parallelization strategies on LLM inference.
V-A Input Batching
Since most of the operators in the decode phase are memory-bound, batching multiple queries can help in improving the decode throughput. Fig. 6 shows how the $\mu_{thr}$ and TPOT increase with increase in batch size. Running a LLaMA2-7B model at int8 precision on NVIDIA’s A100 (80GB) GPU, with 2k input tokens, as we increase the batch from 1 to 44, we see a 10.9$\times$ improvement in the $\mu_{thr}$ while TPOT only increasing by 1.83$\times$. The $\mu_{thr}$ becomes constant after B = 38. After this point, the decode latency increases linearly with batch size. The saturated throughput is 14.5 times the B=1 throughput with only 2.63$\times$ increase in latency. Thus, batching inputs during the decoding phase proves to be highly advantageous.
Figure 6: Impact of batching on decode throughput and TPOT.
Static batching[15, 25] and continuous batching[130, 95, 76, 43] represent the two primary batching mechanisms commonly employed for LLM serving. In static batching, all available requests are batched together, but all the forward passes for these requests are completed before any other requests are run. A key challenge with static batching is higher queuing latency for new requests. For continuous batching, new queries can be added after each forward pass. The prefill stage of the new queries can be run along with the decode phase of the older queries. Thus, it helps reduce the queuing time for new queries but can result in higher tail latency for initial queries. In this work, we gain our insights primarily from static batching, but these insights would work for platforms running continuous batching as well.
V-B Parallelization Strategies
Figure 7: Various parallelization strategies. Each grey box represents an accelerator NPUs and the colored blocks represent the layers. We illustrate how the LLM runs on different NPUs for each strategy. (a) In full TP weights of all layers are equally distributed among all NPUs. (b) With hybrid TP and PP, the layers are distributed among groups of accelerators. Within a group, the layer weights are distributed in tensor parallel fashion. (c) Full PP splits all layers across the NPUs. In all cases, once the model has been partitioned via TP/PP, DP is employed across the remaining NPUs in the platform.
Meeting SLO demands another reason to scale up the number of accelerators in the platform. However, there is a lack of clarity about the impact of the parallelism strategy. An optimal parallelization strategy may vary depending on the model and workload configuration. We also observe drastic contrast in how different parallelization strategies work during the prefill and decode stage. Fig. 7 shows the example of how model weights and KV cache split with different degrees of TP + PP. We restrict our evaluations to TP and PP in this work across the diverse LLMs. We leave platform implications of EP and SP for MoEs as future work.
Fig. 8 compares different LLM models on HGX: H100x8. We use a workload with $S_{b}=4,\tau_{p}=4000,\tau_{d}=256$.
Key takeaways for choosing parallelization strategy: Prefill: PP can serve many more queries than TP. PP=8 can 2.32x, 1.38x more query compared to TP=8 for Mixtral 8x7B and GPT3-175B. Decode: TP only or Mixture of TP + PP is generally is superior for throughput. On the other hand, higher PP enables the platform to accommodate much higher batch sizes. For example, HGX:H100x8 with PP=8 can accommodate GPT3-175B, which can accommodate a batch size of 256 due to micro-batching, while TP gives an OOM error after B=32.

### v2

@@ II-C Metrics for LLM Serving
III GenZ : Generative LLM Analyzer
GenZ is an analytical framework that can be used to study different LLM model architectures combined with the latest software optimizations on distributed current and next-generation NPU platforms. GenZ has three key components: 1) model profiler, 2) NPU characterizer, and 3) platform characterizer. We show an overview of GenZ in Fig. 2 and discuss each component in the following sections.
Real Life
Application
@@ III GenZ : Generative LLM Analyzer
Beam
Width
TTFT / TPOT
(s) / (ms)
Question Answering 1000/200 4 0.2 / 10
Chat Services 3000/1000 2 0.2 / 10
QA + RAG 10000/200 4 0.4 / 10
Text Summarization 15000/1000 4 2 / 20
Code Generation 20000/50 4 0.5 / 20
TABLE III: Representative Use Cases of LLM models and their representative input hyperparameters.
Model $D_{model}$ Layers
Heads
($H_{kv}$)
$W_{ff}$
Experts
Gemma2-2B [51] 2304 26 8(4) 4 -
LLaMA3-8B [52] 4096 32 32(8) 3.5 -
Gemma2-27B [51] 4608 46 32(16) 8 -
Mixtral-8x22B [47] 6144 56 48(8) 2.66 8(2)
LLaMA3-70B [52] 8192 80 64(8) 3.5 -
GPT3-175B [7] 12288 96 96 4 -
LLaMA3-405B [52] 16384 126 128(8) 3.25 -
GPT4-1.8T [49] 10752 120 84 4 16(2)
Dense-5T 49152 128 192(24) 4 -
MoE-10T 13824 128 108(12) 4 32(4)
TABLE IV: Different LLM architectures evaluated in this work.
III-A Model Profiler
@@
Table IV shows the parameters of various LLM models that we study in this work. These models serve as a representative set of current and future LLMs. GenZ model can model different LLM architectures, including MoE-based LLMs (Mixtral-8x7B [46] and GPT4) and Mamba-based models [58]. LLaMA2, LLaMA3, Mixtral, and GPT3 architecture are available openly. We represent GPT4 as a 1.8T parameter mixture-of-experts model, with 120 layers [49, 50, 17]. A single layer of GPT4 has 16 experts of  [^1] B parameters each, and two experts are activated for each token. We also hypothesize two future LLM models  [^2] We do not train these models for accuracy, which is not the focus of this work. We use it to study its computational behavior. to represent scaled-up Dense-GQA and MoE models.
@@
GenZ’s support for different architectures enables it to model all the SOTA LLMs. For each new model, GenZ uses the huggingface AutoModels [59] to determine the exact shape of each operator in the model. We store the operator dimension of layers. Using these stored variables, we calculate the number of operations, operator execution engine, operator residency information, operator size, KV cache estimation, collective sizes, and collective groups.
@@
Saving model operators offline allows us to profile larger context lengths for any LLM model quickly. For example, LLaMA2-7B has a context length limit of 4K, but we can extend the context length to any size using this offline modeling.
@@
We also model popular optimizations in this stage of model profiling, including Kernel Fusion (Flash Attention [33, 60, 61], Segment KV Caching [62]), model quantization, chunked prefilling [35, 34], speculative decoding [63], and beam search [55, 56]. Since this work focuses on LLM inference, we use FP8 model quantization for all our experiments and results unless specified otherwise.
III-B NPU Characterizer
Our smallest hardware unit is the accelerator (alternatively referred as the NPU), as shown in Fig. 2. We assume each NPU has a certain number of compute cores, which can perform $FLOPS$ number of operations per second. We use a variable ($Eff_{C}$) to account for the inefficiencies caused by the software and memory synchronization issues. For modeling real systems, GenZ uses the runtime of real systems (e.g. time to execute matmul operation on A100 GPU). This execution time is used to calculate the efficiency factor. This is the same methodology as adopted by previous works like Vidur [15]. For modeling future hypothetical NPUs, the microarchitecture of NPUs plays a crucial role in determining runtime of each operator. To effectively model the effects of microarchitecture, GenZ leverages external high-fidelity tools, such as Timeloop [64] and SCALE-sim  [41], which specialize in simulating individual NPU dataflows and microarchitecture details. GenZ generates the operator dimensions for a given model architecture, system optimization, model optimization and parallelism strategy. This operator dimension is feed to the external tool to get the operator runtime, and thus get the hardware efficiency factor. Section VII-D shows a case study in which we model different microarchitecture using SCALE-sim[41] and study their effect on the prefill stage of LLama-3-8B.
Each NPU provides access to two external memories (fast and slow). Faster (smaller) memory represents an HBM/DDR Memory Bank providing high BW($BW_{mem}$), while Slower (Larger) memory could be PCIe-accessible CPU or CXL-accessible SSD/Flash ($BW_{omem}$) for offload. We also use an efficiency factor($Eff_{mem},Eff_{omem}$) with the memory link for accurate memory access time.
Since all operations in LLM inference have pre-determined shapes and we use all model weights and KV cache uniformly, a smart compiler can try to maximize the overlap between the operator computation and memory fetches. Thus, we analyze the model’s performance on an operator-by-operator basis. For each operator, we calculate its corresponding operations per sec ($C_{op}$) and the number of memory accesses($M_{op}$). We follow a roofline-based approach combined with separate efficiency factors (extracted from real hardware or open source simulators) for computation FLOPs and memory BW to calculate each operator’s runtime on the accelerator.
$T_{op}$ $=max(\frac{C_{op}}{FLOPS\times Eff_{C}},\frac{M_{op}}{BW_{mem}\times Eff_{mem}})$ (1)
This simple yet effective modeling methodology is perfect for quickly estimating LLM serving trends on different hardware. In Section  III-D1, we show that using these efficiency factors, we can simulate the trends generated on real GPUs.
III-C Platform Characterizer
One of the key features of GenZ is its ability to simulate multi-dimensional network topologies for LLM inference. GenZ defines an inference platform as multiple NPUs connected through a multi-dimensional interconnection network (ICN), for scale-up and scale-out, as shown in Fig. 2. Each dimension in ICN has the following properties: link latency ($T_{link}$), network link bandwidth ($BW_{link}$), and network link efficiency ($Eff_{link}$).
Figure 4: Various parallelization strategies for neural network training and inference. Each colored box represents an accelerator (NPU), and the numbers correspond to model layers.
Parallelism. Typically distributed LLMs are served using five different types of parallelism strategies: Data Parallel (DP), Tensor parallel (TP)  [65], Pipeline parallel (PP)  [66], Expert parallel (EP) [45] (Only for MoE models) and Sequence parallel (SP) [67, 68] (primarily for training long sequence). Fig. 4 shows the example of how different parallelism splits the input tokens and model weights. GenZ handles the overlap of physical topology and logical parallelism topology. For our experiments, we use the parallelism order as TP:EP:PP. This order points to how the NPUs are physically located to one other. For TP:EP:PP, the NPUs doing TP are physically the closest; next, NPUs with EP, and finally, nodes with PP. The degree of parallelism can be arbitrary, and GenZ will correctly map the logical parallelism mapping on the multi-dimension physical network topology defined by the user.
Collectives. For each degree of parallelism, GenZ generates the required collective pattern(s). There can be five kinds of collective patterns: AllReduce (TP & EP), All-to-All (EP), Send-Recv (PP), AllGather (SP & TP), and ReduceScatter (TP). GenZ allows the all-reduce collective to be broken down into ReduceScatter followed by AllGather for hiding the communication latencies. To get the runtime for each collective, we simulate it by calling the ASTRA-sim [69] system-layer as that provides implementations for diverse topology-aware collective algorithms. In our modeling, we also have a knob to decide whether to overlap collectives with compute tasks or execute them sequentially. For this work, we use non-overlapping communication similar to SOTA frameworks [70].
Figure 5: GenZ (a-d) LLM inference modeling against vLLM inference in prefill and decode stage for varying batch size on various platforms with different $\tau_{p}$. Line color blue, red, and green represent LLaMA2-7B, LLaMA2-13B, and OPT-175B.
III-D GenZ Implementation and Validation
III-D1 GenZ Implementation and Runtime
GenZ is implemented in over 5,000 lines of Python code and packaged as a single Python module, enabling seamless integration into design space exploration (DSE) workflows via a simple pip installation. The tool is computationally lightweight, requiring approximately 30 milliseconds per forward pass on a standard 8-core CPU, given a batch size of 64 and 1024/256 input/output tokens. Runtime efficiency is further enhanced through operator reuse: GenZ identifies and skips redundant computations by sharing runtime estimates across layers.
III-D2 GenZ Validation
@@
We validate GenZ on five different real hardware platforms: a) HGX [71] box with 8 NVIDIA H100 SXM GPUs (80 GB) fully connected by NVLinks, b) 2xA100, c) Intel Gaudi2, d) AMD MI300x, e) 8xSambanova SN40L. To assess GenZ ’s accelerator modeling, we analyze prefill, decode, and chunking trends across these systems.
Efficiency Factors: Our measured efficiency factors are derived from profiling real NPUs, following a methodology similar to Vidur [15]. We execute the same kernel multiple times and measure average utilization to obtain realistic efficiency estimates. LLM inference frameworks like vLLM execute static PyTorch computational graphs, ensuring consistent hardware utilization across runs. We validate our approach by comparing predicted runtimes against median measured runtimes, minimizing the impact of outliers. While we observe linearity in the validation data, it arises naturally from the predictable scaling of LLM inference workloads, where execution time for key bottlenecks (e.g., matrix multiplications, attention) scales proportionally with input size or batch size, rather than from manual tuning, as also observed in other works [72]. For each validation study, we report the efficiency factor for the system used.
Prefill & Decode validation: Using vLLM, we evaluate LLaMA2-7B, LLaMA2-13B, and OPT-175B across various platforms, using randomly generated dummy paragraphs as input, ranging from 500 to 2000 tokens, with each model generating a fixed 200 output tokens. Fig. 5(a)-(d) compares the models’ prefill TTFT and decode throughput on NVIDIA platform. The geomean error in prefill and decode predictions between real and GenZ-predicted values is 2.73% and 1.85%, respectively, across different models and platforms. The average empirically-measured efficiency factors used for different hardware configurations are V100: 0.45, A100: 0.4, 1×H100: 0.55, 2×H100: 0.64, 4×H100: 0.66, and 8×H100: 0.75.
Chunked Validation: We run chunked serving for Llama2-7B (bf16) on 2xA100 using vLLM engine varying batch sizes (1-32), input lengths (512-2048), and chunk sizes (256,768). Fig. 6 compares actual end-to-end serving times against GenZ estimates, yielding a geomean error of 1.43%.
Validation across architectures: We also validate GenZ against three other popular architectures [^3] We were unable to get access to the physical node for these architecture, so we used the number from LLM-Inference-Bench [72]. The raw data was accessed from https://github.com/argonne-lcf/LLM-Inference-Bench: (i) 8xSambanova SN40L [73], (ii) 1xAMD MI300X running vLLM [32] and (iii) 1xIntel Gaudiv2 running deepspeed [34]. Fig. 7 compares the request serving time on these platforms when running LLaMA3-8B (bf16) with batch size 16, varying input/output lengths, GenZ achieves a geomean error of 5.82% across all different architectures.
Figure 6: Validation results comparing GenZ against vLLM inference running chunked prefill on 2xA100 (Eff=0.35). Llama2-7B with varying batch sizes, input lengths, and chunk sizes.
Figure 7: Validating different architectures running Llama3-8B (bf16). SN40L uses Sambaflow framework (Eff=0.9), MI300X uses vLLM (Eff=0.25) and Gaudi2 uses deepspeed (Eff=0.6).
@@
Platform interconnect validation: To verify the collectives time at the platform scale, we benchmark all-reduce communication primitive with nccl-tests [74], a communication primitive performance benchmark for NVIDIA GPUs. GenZ ’s collective communication times are validated with a platform size of 2 GPUs, 4 GPUs, and 8 GPUs. For collective communications, we observed an average link efficiency of 75% for NVLINK, which gives an effective link BW of 350 GB/s for each GPU in HGX box. We profile all model (Table IV for different input and output lengths and collect their all-reduce (AR) message sizes. The message size of each AR call is very small ($<$ 128 KB) for decode, while for prefill, per call message size is a few hundred MBs. Fig. 8 compares the real hardware latency for three different platform sizes and the corresponding latency generated by GenZ . Since the prefill and decode stages have stark differences in message size, we verify the collective time for prefill and decode independently. We found that for decode message sizes, the link latency, $T_{Link}$ is the dominating, thus the latency seems almost constant. While prefill AR messages are large enough, that the link bandwidth, $BW_{Link}$, is the main contributor to collective time. Collective for all datapoints, there is a 3.89% and 2.7% geomean error for decode and prefill message sizes between real values and GenZ values.
@@
These results confirm that GenZ accurately captures inference trends observed on real hardware.
IV Impact of Inference Optimizations on Hardware
LLM inference is one of the most active fields of research in recent years. This has led to a rapid rise in the number of innovations and optimizations being introduced. Table V summarizes a few of the most popular techniques and their impact on compute and memory requirements from the AI inference platform. These fall into broadly 3 buckets, i.e., 1 Model architecture change, 2 System and algorithmic optimizations which don’t change the model quality, and finally 3 Algorithmic optimization with model quality changes.
GenZ supports almost all of the techniques shown in Table V, with being the only framework that supports modeling MoEs, mamba, and hybrid models, at the same time supporting sliding window attention, speculative decoding, weight sparsity, KV pruning, and mixed precision computation and storage.
Technique Comment
Compute / Memory
Impact
1 Foundational model architecture change (Requires pre-training)
MQA/GQA Fewer KV heads - /
MoE Sparesely activated FFNs /
Sliding Window Smaller attention window /
Layer-wise KV sharing Multiple layers share KV cache - /
2 Lossless System optimization without any impact on model quality
Flash Attention Reduced memory accesses - /
Chunking prefill Prefill Split + Decode /
Parallelism Distributed inference - /
Speculative Decoding Draft model predicts tokens /
3 Lossy Model optimization with impact on model quality
Quanitization Reduced bitwidths /
Weight Sparsity Removing weights - /
KV pruning Removing KV cache tokens /
Mixed precision
Different bitwidth for
storage and computation
TABLE V: Various techniques for optimizing LLMs and their impact on compute and memory requirements.
In this section, we use GenZ to model the impact of three optimization techniques and provide insights to build next-generation AI platforms for running with those techniques.
IV-A Chunked Prefill
Chunked prefill [35] or chunking [53, 54] or SplitFuse [34] combines the compute-bound prefill and memory-bound decode stages of LLM generation to provide better throughput. All outstanding decode batches are processed parallelly in the chunk. Additional tokens are padded by outstanding prefill requests’ tokens to construct chunks of fixed size. This ensures that each forward pass always has a fixed number of tokens. Since the number of tokens in the forward pass is fixed, the runtime of most layers is always fixed.
Figure 9: Runtime breakdown of inference latency with varying chunk size and decode batches.
To understand the effect of chunking, we run two models, i.e., GPT-3 and LLama-3.1-405B, with tensor parallelism of 4 on GB200-like NPU. With $\tau_{p}=4096,\tau_{d}=1024$. We vary the decode batch size from 1 to 128 and the chunk size from 256 to 2048. Fig. 9 shows the runtime breakdown of this study. GPT-3 is unable to fit larger batch sizes as the model has dense architecture. LLama-405B, even with a larger model size, fits a much higher batch size as it uses GQA architecture. For a given chunk size, we observe that the linear GEMM layers have nearly constant latency. Only the latency of logit and attend layers (Q.K’ + Softmax + S.V) increases. This is due to the fact that these layers are also memory-bound irrespective of the context length and the batch size. In the GPT-3 dense model, as decoded batches accumulate, the growing KV cache (red bar components) becomes a bottleneck, thus memory bandwidth becomes the bottleneck for dense models. For the LLama-405B model with GQA architecture, the multi-headed attention(logit +attend) component is a very small part of the overall latency, and thus, latency remains didn’t increase significantly with decode batch size.
$\bullet$ Memory bottleneck for dense models: Memory BW would remain a significant bottleneck due to larger KV caches. A larger memory size would also be required to process more decode batches in parallel. $\bullet$ Compute bottleneck for GQA models: Memory critical layers contribute a very small portion of the total runtime. The compute-bound layers are the primary bottleneck for running models with GQA when doing chunking.
IV-B Speculative Decoding
Speculative decoding(SD)  [63, 28, 75, 76, 77, 78] is a system level optimization technique for accelerating token generation without compromising accuracy. It uses a smaller draft model to generate multiple speculative token sequences in an auto-regressive fashion. These tokens serve as ”guesses” for the large target model. The large target model evaluates these guesses in a single pass. It either accepts the suggested tokens if they align with its probability distribution or rejects them and adjusts the next output accordingly. On the flip side, if any token generated by the draft model is rejected by the target model, all subsequent tokens are dropped. Thus, the overall output throughput would be reduced.
We use two hyperparameters for modeling SD in GenZ : i) N = number of tokens generated by the draft model before the large full model checks them. ii) $\gamma$ = probability of each token generated by the smaller model being accepted by the target model. This helps us estimate the expected tokens per pass of the target model as
$E[T]=\sum_{i=1}^{N-1}i\cdot\gamma^{i}\cdot(1-\gamma)+N\cdot\gamma^{N}$
[…]
Figure 10: Comparing baseline LLM inference against inference using speculative decode execution with a draft model and a number of parallel decode tokens, $N=5$.
Fig. 11 shows decode throughput on GB200-like NPU with TP=2. We test with $N\in\{4,16\}$ $\gamma\in\{0.7,0.9\}$ for Llama-3.1-70B (draft model: Llama-3.1-8B) and Gemma-2-27B (draft model: Gemma-2-2B). We also vary the $\tau_{p},\tau_{d}\in\{512,1024,2048\}$. The dashed lines represent the model throughput without SD. 1 We see that as the number of parallel tokens, N, increases, the throughput goes down. This is because of the increased cost of running the draft model. For $\gamma=0.7$ N=16, both the models perform worse than they would without SD. 2 For a given number of parallel decode tokens, i.e., fixed N, $\gamma$ is directly linked to the throughput gains. We see that for N = 4, $\gamma=0.5$ has roughly the same throughput as the model w/o SD.
The draft models, Gemma-2-2B and Llama-3.1-8B, require 10.8% and 9.6% extra memory for weights and 40% and 28% extra memory for KV cache. For a batch of 4 requests with input/output length 32K, the total additional memory requirements would be 24.7% and 28.2%, respectively.
$\bullet$ Larger memory capacity: Keeping 2 models on-device with their corresponding KV cache requires HW with larger memory capacity compared to running a single model. $\bullet$ Compute bottleneck: With multiple parallel tokens being fed to the target model, most operators can be pushed to the compute-bound region. This means that more layers are compute-bound than memory-bound.
Figure 11: Decoding Throughput with speculative decoding.
IV-C Optimal Parallelization Strategies for MoEs
While tensor parallelism is generally known to be the best parallelism type for dense LLMs during inference  [72], it is still unclear which parallelism technique would work well with MoE models. Using GenZ , we explore different parallelism strategies for a popular MoE, Mixtral-8x22B, on an H100x8 node connected by a switch network shown in Fig. 12. We assume that tokens are distributed among the experts in a balanced fashion for prefill.  [^4] Exploring the effect of load imbalance among experts is left as future work. However, the small number of tokens makes it very unpredictable during the decode. Thus, TPOT of Mixtral 8x22B on 4 H100 with expert parallelism can vary between 3.23 ms (All tokens distributed equally) and 11.33 ms (All tokens going to a single expert) for batch 32. Although different scenarios and assumptions could result in choosing different parallelism as optimal, we believe our tool, GenZ can be used to find optimal parallelism for future MoEs on any HW platforms.
$\bullet$ EP is preferred when all experts activated: For prefill and chunked stages, load balanced expert activation, EP is the best parallelism strategy. In case there is an imbalance among experts, EP could perform significantly worse. $\bullet$ Mix of TP and EP is preferred when partially activated experts: For the decode stage, where only a subset of experts are activated with very few tokens being routed to each expert, TP only or TP + EP is generally superior for throughput.